OpenStudy (anonymous):

find the length of the curve y=sinx from x=0 to x=pi/2

zepdrix (zepdrix):

Arc length,\[\Large\bf\sf L\quad=\quad \int\limits ds\]

zepdrix (zepdrix):

Where ds is an immeasurably small piece of arc. We add up all of the pieces of arc.

zepdrix (zepdrix):

\[\Large\bf\sf ds\quad=\quad \sqrt{1+\left(\frac{dy}{dx}\right)^2}dx\]